OpenStudy (anonymous):

find instantaneous rate of change for function f(x)= log x, when x = 2.

OpenStudy (anonymous):

the answer is not 1/2.

OpenStudy (ash2326):

what's the answer given

OpenStudy (ash2326):

is it log to the base 10 or e

OpenStudy (anonymous):

log base 10

OpenStudy (anonymous):

?

OpenStudy (ash2326):

then change base first , we know that \[\log_{a} x= \log_{e} x/\log_{e} a\] so here \[\log_{10} x=\log_{e} x/\log_{e} 10\] so \[f(x)=\log_{e} x/\log_{e} 10\] now find f'(x) which is \[f'(x)=1/(x*\log_{e} 10)\] x=2 so \[f'(2)=1/(2*\log_{e} 10)\] \[\log_{e} 10=2.302\] \[f'(2)=1/(2*2.302)\] \[f'(2)=0.2172\]